# Roadmap for electronic structure, anharmonicity, and electron-phonon calculations in locally disordered inorganic and hybrid halide perovskites

### v1

@@ II.4 Anharmonic phonons via the ASDM
where $F$ is the true free energy, $F_{\rm vib}$ is the vibrational free energy of the reference system, and $U$ is the PES of the actual system with the thermal average .$\braket{.}_{T}$ taken with respect to the eigenstates of the reference system. The inequality in Eq. (7) suggests that variational minimization of the trial free energy $F_{0}(T)$ of the simpler reference system provides an upper bound on the true free energy $F(T)$ of the more complex anharmonic system.
@@ III.1 Methodology
    When multiple ZG reference or polymorphous structures are used to compute a property $O$, perform a Boltzmann-weighted average, defined as:
    $\braket{O}_{j}=\sum_{j}\frac{1}{Z}e^{-F_{j}/k_{\rm B}T}O_{j},$ (11)
    where .$\braket{.}_{j}$ is a configurational average, $F_{j}$ is the free energy of the structure $j$ [Eq. (7)], and $Z=\sum_{j}e^{-F_{j}/k_{\rm B}T}$ is the partition function. When no vibrational contributions are considered, the free energy is replaced by the DFT KS total energy.
The choice to first obtain a reference structure, i.e. relaxing the organic molecules before relaxing the entire halide perovskite system, rather than using the monomorphous structure, helps prevent artifacts arising from the net dipole moments of the molecules, particularly in the case of MA-based compounds. We note that for inorganic halide perovskites the reference structure is identical to the monomorphous structure. The generation of a reference structure in hybrid systems is also rationalized by the mostly distinct time scales and dynamics of the organic and inorganic components. The organic molecules exhibit, on majority, faster internal vibrations (energy scales $>$ 100 meV) and slower rotational and librational motions (energy scales $<$ 4 meV) compared to the inorganic lattice, whose phonon dynamics typically occur at intermediate energy scales of 4 – 35 meV Zacharias et al. (2023b). Starting from random molecular orientations allows the molecules to explore their local potential energy landscape, where their dipoles adjust to the electrostatic environment created by the interplay with the fixed inorganic lattice. Relaxing the molecules first ensures that these slower rotational and librational modes settle into stable local minima without interference from the lattice dynamics. Additionally, it allows to mimic the effects of experimentally observed slow and stochastic molecular librational relaxations Even et al. (2016), which can not be deduced from the unstable lattice modes of the monomorphous structure using Eq. (6). In the subsequent step, the whole structure is relaxed together, enabling the inorganic lattice to adjust and couple with the pre-optimized molecular orientations. This two-step approach captures the polymorphous nature of the cubic phase, where local distortions and dynamic interactions between the organic and inorganic sublattices coexist, in a quasistatic manner. Additionally, it provides a reference structure that serves as a well-defined baseline for subsequent phonon quasiparticle calculations, such as those performed within the framework of the SCP theory.
@@
Figure 4: Structures considered throughout this study. Reference structures can contain symmetrized-constrained or relaxed random molecular orientations. These structures lead to distinct polymorphous structures after geometry optimization. The ZG structures, shown as green, incorporate electron-phonon coupling effects at finite temperatures ($T$).
@@ III.2 Polymorphous structures, total energy lowering, and electronic structure correlations
We first apply steps 1-4 in the main methodology for generating 10 different reference and their corresponding polymorphous structures for 12 materials. All calculations refer to $2\times 2\times 2$ supercells using the PBEsol functional; more computational details are available in Sec. VI. Figure 2(b) presents the Boltzmann-weighted average [Eq. (11)] of the total energy lowering ($\Delta U$) calculated for the polymorphous structures relative to the reference structures, assuming a temperature $T=300$ K. We note that testing different temperatures from 30 to 1000 K affects the values by no more than 4 meV per formula unit. On average, we find that the potential well depth, $\Delta U$, is shallower in FA-based compounds compared to MA-based compounds, which in turn exhibit a shallower potential well than Cs-based compounds. The observed trend arises from differences in correlated local disorder in the inorganic network, where, FA-based compounds exhibit the least structural distortions, MA-based compounds show moderate disorder, and Cs-based compounds experience the highest degree of octahedral tilting and bond length fluctuations, leading to a deeper potential well. This behavior is mainly explained by the size of the A-site cation. Larger organic molecules, such as FA, create a more sterically constrained environment that limits structural flexibility, whereas smaller cations like MA and Cs permit greater lattice distortions and allow local disorder to emerge more readily. This trend is consistent with the tolerance factor Zacharias et al. (2025); Kieslich et al. (2014); Filip and Giustino (2018); Li et al. (2022), where lower values, associated with smaller A-site cations, favor increased positional polymorphism.
@@
We quantify the degree of local disorder in the inorganic network by calculating the average B-X bond length ($\Delta l_{\rm B-X}$) and B-X-B bond angle variations ($\Delta\theta_{\rm B-X-B}$) relative to the idealized reference structures and demonstrate their correlations with the induced band gap increase ($\Delta E_{\rm g}$) and hole/electron effective mass enhancements ($\Delta m^{*}_{\rm h}$/$\Delta m^{*}_{\rm e}$), as shown in Figs. 2(c)-(h). In all cases we obtain strong correlations indicated by the Pearson correlation, r, and R2 coefficients, ranging from 0.84 to 0.96. Consistent with our findings on the potential well depth [Fig. 2(a)], the data categorize into three distinct groups. FA-based compounds, lying on the left side of the plot, exhibit $\Delta l_{\rm B-X}$ between 0.013 – 0.024 Å  and $\Delta\theta_{\rm B-X-B}$ ranging from 8.2 – 13.0. MA-based compounds, located in the middle, show $\Delta l_{\rm B-X}$ from 0.027 – 0.060 Å  and $\Delta\theta_{\rm B-X-B}$ between 12.1 – 20.2. On the right, Cs-based compounds display the largest structural deviations, with $\Delta l_{\rm B-X}$ spanning 0.057 – 0.085 Å  and $\Delta\theta_{\rm B-X-B}$ ranging from 21.1 – 26.5. Notably, MA-based compounds cover the widest range of structural variations, which we attribute to the larger dipole moments of the MA cations that can significantly influence the electronic structure Quarti et al. (2015); Motta et al. (2015).
@@
Figure 5: Dependence of the average B-X-B bond angle decrease on the average B-X bond length variation. The 120 data points correspond to 10 polymorphous configurations generated for 4 FA (black), 4 MA (red), and 4 Cs (blue) -based compounds. The straight thick line represent the least squares regression fit to the data, with the slope indicated on the plot. The shaded region represents three standard deviations on either side of the lines.
@@
[…]
@@ III.3 Phonon anharmonicity and diffuse scattering
[…]
@@ III.4 Free energies, temperature-dependent band gaps, and effective masses
We remark that the 10 polymorphous structures explored for each cubic hybrid halide perovskite in 2$\times$2$\times$2 supercells yield distinct DFT total energy lowerings and phonon spectra. These variations arise from the sensitivity of local disorder to molecular orientations as discussed previously. To account for the influence of local disorder and thermal effects on the electronic structure across various ZG polymorphous configurations, we compute the free energy of each configuration and use it to obtain Boltzmann-weighted averages [Eq. (11)]. Figures 7(a)-(f) show calculations for the relative free energies of each ZG polymorphous configuration and Boltzmann averaged band gaps and effective masses of cubic MAPbBr3 and FAPbI3. The free energies in Figs. 7(a,b) are evaluated with respect to the least stable configurations, r2 and r5. We also include the relative free energy calculated for the ZG polymorphous configuration obtained by combining the ASDM with geometry optimization for the MSS structure. Although starting from the MSS structure yields one of the most stable configurations, it does not lead to the absolute minimum in stability. Other random molecular orientations facilitate local disorder within 2$\times$2$\times$2 supercells more effectively, leading to lower energy ZG polymorphous structures.
@@
In Figs. 7(d) and (f), we present the temperature dependence of the absolute values of hole and electron effective masses in cubic MAPbBr3 and FAPbI3, exhibiting a linear relationship with temperature. Compared to calculations for the reference structures, the effective masses obtained using ZG polymorphous configurations are almost twice as large. In fact, for the ZG reference structure of cubic MAPbBr3 we obtain $m^{*}_{\rm h,ref}=0.15$, $m^{*}_{\rm e,ref}=0.14$, and a reduced effective masses of $\mu_{\rm ref}=0.07$ at $T=300$ K, where $m_{0}$ is the free electron mass. In contrast, for the ZG polymorphous structure of cubic MAPbBr3, these values increase substantially to $m^{*}_{\rm h,poly}=0.26$, $m^{*}_{\rm e,poly}=0.23$, and $\mu_{\rm poly}=0.12$ at the same temperature. The corresponding values at $T=300$ K for cubic FAPbI3 are $m^{*}_{\rm h,ref}=0.10$, $m^{*}_{\rm e,ref}=0.08$, $\mu_{\rm ref}=0.04$, $m^{*}_{\rm h,poly}=0.16$, $m^{*}_{\rm e,poly}=0.13$, and $\mu_{\rm poly}=0.07$. We note that our calculations are at the DFT-level and can be further improved by considering quasiparticle corrections within the $GW$ approximation, which are known to enhance the effective masses Huang and Lambrecht (2013); Umari et al. (2014); Brivio et al. (2014); Bokdam et al. (2016); Mosconi et al. (2016); Davies et al. (2018); Wang et al. (2020b); Filip et al. (2021); Sajedi et al. (2022) In Sec. IV, we provide a more detailed discussion how the reduced effective masses of polymorphous structures compare to experimental measurements and $GW$ calculations.
Table 1: Lattice constants ($L$), DFT band gaps ($E_{\rm g}$), PBE0 band gaps ($E^{\rm PBE0}_{\rm g}$), DFT hole ($m^{*}_{\rm h}$) and electron ($m^{*}_{\rm e}$) effective masses, and reduced effective masses ($\mu$) of the reference (r) and polymorphous (p) structures of inorganic and hybrid halide perovskites. Band gaps and effective masses are calculated using $2\times 2\times 2$ supercells (40 and 96 atoms), $3\times 3\times 3$ uniform ${\bf k}$-grids, and a plane wave cutoff of 120 Ry, including the effect of spin-orbit coupling. DFT corrections to the band gaps [$\Delta E_{\rm g}$($T$)] and effective masses [$\Delta m^{*}_{\rm h}$($T$)] due to electron-phonon coupling and thermal expansion at temperatures $T$ are calculated with the ASDM using $4\times 4\times 4$ ZG supercells (320 and 768 atoms), $1\times 1\times 1$ uniform ${\bf k}$-grids, and a plane wave cutoff of 60 Ry, including the effect of spin-orbit coupling. Values were obtained as averages over 10 distinct configurations of the reference and polymorphous structures. Experimental band gaps of the cubic phases, $E^{\rm expt.}_{\rm g}$, and experimental reduced effective masses of the low-temperature phases ($\mu^{\rm expt.}$) are reported for comparison. $\Delta_{s}$ represents an empirical correction to the calculated band gaps to match $E^{\rm expt.}_{\rm g}$, which is consistent with the PBE0 correction $E^{\rm PBE0}_{\rm g}-E_{\rm g}$.

### v2

@@ II.4 Anharmonic phonons via the ASDM
where $F$ is the true free energy, $F_{\rm vib}$ is the vibrational free energy of the reference system, and $U$ is the PES of the actual system with the thermal average $\braket{.}_{T}$ taken with respect to the eigenstates of the reference system. The inequality in Eq. (7) suggests that variational minimization of the trial free energy $F_{0}(T)$ of the simpler reference system provides an upper bound on the true free energy $F(T)$ of the more complex anharmonic system.
@@ III.1 Methodology
    When configurational entropy is accounted for, i.e. multiple ZG reference or polymorphous structures are used to compute a property $O$, perform a Boltzmann-weighted average, defined as:
    $\braket{O}_{j}=\frac{1}{Z}\sum_{j}e^{-F_{j}/k_{\rm B}T}O_{j},$ (11)
    where $\braket{.}_{j}$ is a configurational average, $F_{j}$ is the free energy of the structure $j$ [Eq. (7)], and $Z=\sum_{j}e^{-F_{j}/k_{\rm B}T}$ is the partition function. When no vibrational contributions are considered, the free energy is replaced by the DFT KS total energy.
The choice to first obtain a reference structure, i.e. relaxing the organic molecules before relaxing the entire halide perovskite system, rather than using the monomorphous structure, helps prevent artifacts arising from the net dipole moments of the molecules, particularly in the case of MA-based compounds. We note that for inorganic halide perovskites the reference structure is identical to the monomorphous structure. The generation of a reference structure in hybrid systems is also rationalized by the mostly distinct time scales and dynamics of the organic and inorganic components. The organic molecules exhibit, on majority, faster internal vibrations (energy scales >> 100 meV) and slower rotational and librational motions (energy scales << 4 meV) compared to the inorganic lattice, whose phonon dynamics typically occur at intermediate energy scales of 4 – 35 meV Zacharias et al. (2023b). Starting from random molecular orientations allows the molecules to explore their local potential energy landscape, where their dipoles adjust to the electrostatic environment created by the interplay with the fixed inorganic lattice. Relaxing the molecules first ensures that these slower rotational and librational modes settle into stable local minima without interference from the lattice dynamics. Additionally, it allows to mimic the effects of experimentally observed slow and stochastic molecular librational relaxations Even et al. (2016), which can not be deduced from the unstable lattice modes of the monomorphous structure using Eq. (6). In the subsequent step, the whole structure is relaxed together, enabling the inorganic lattice to adjust and couple with the pre-optimized molecular orientations. This two-step approach captures the polymorphous nature of the cubic phase, where local distortions and dynamic interactions between the organic and inorganic sublattices coexist, in a quasistatic manner. Additionally, it provides a reference structure that serves as a well-defined baseline for subsequent phonon quasiparticle calculations, such as those performed within the framework of the SCP theory.
@@
Figure 4: Structures considered throughout this study using the example of MAPbBr3. Reference structures can contain symmetrized-constrained or relaxed random molecular orientations. These structures lead to distinct polymorphous structures after geometry optimization. The ZG structures, labelled as green, incorporate electron-phonon coupling effects at finite temperatures ($T$). For the ZG structures, we considered a $\times4$$\times$4 supercell (768 atoms) to accommodate a broader sampling of vibrational modes.
@@ III.2 Polymorphous structures, total energy lowering, and electronic structure correlations
We first apply steps 1-4 in the main methodology for generating 10 different reference and their corresponding polymorphous structures to account for configurational entropy effects in 12 materials. All calculations refer to $2\times 2\times 2$ supercells using the PBEsol functional; more computational details are available in Sec. VI. Figure 2(b) presents the Boltzmann-weighted average [Eq. (11)] of the total energy lowering ($\Delta U$) calculated for the polymorphous structures relative to the reference structures, assuming a temperature $T=300$ K. We note that testing different temperatures from 30 to 1000 K affects the values by no more than 4 meV per formula unit. On average, we find that the potential well depth, $\Delta U$, is shallower in FA-based compounds compared to MA-based compounds, which in turn exhibit a shallower potential well than Cs-based compounds. The observed trend arises from differences in correlated local disorder in the inorganic network, where, FA-based compounds exhibit the least structural distortions, MA-based compounds show moderate disorder, and Cs-based compounds experience the highest degree of octahedral tilting and bond length fluctuations, leading to a deeper potential well. This behavior is mainly explained by the size of the A-site cation. Larger organic molecules, such as FA, create a more sterically constrained environment that limits structural flexibility, whereas smaller cations like MA and Cs permit greater lattice distortions and allow local disorder to emerge more readily. This trend is consistent with the tolerance factor Zacharias et al. (2026); Kieslich et al. (2014); Filip and Giustino (2018); Li et al. (2022), where lower values, associated with smaller A-site cations, favor increased positional polymorphism.
@@
We quantify the degree of local disorder in the inorganic network by calculating the average B-X bond length ($\Delta l_{\rm B-X}$) and B-X-B bond angle variations ($\Delta\theta_{\rm B-X-B}$) relative to the idealized reference structures and demonstrate their correlations with the induced band gap increase ($\Delta E_{\rm g}$) and hole/electron effective mass enhancements ($\Delta m^{*}_{\rm h}$$\Delta m^{*}_{\rm e}$), as shown in Figs. 2(c)-(h). In all cases we obtain strong correlations indicated by the Pearson correlation, r, and R2 coefficients, ranging from 0.84 to 0.96. Consistent with our findings on the potential well depth [Fig. 2(a)], the data categorize into three distinct groups. FA-based compounds, lying on the left side of the plot, exhibit $\Delta l_{\rm B-X}$ between 0.013 – 0.024 Å  and $\Delta\theta_{\rm B-X-B}$ ranging from 8.2 – 13.0. MA-based compounds, located in the middle, show $\Delta l_{\rm B-X}$ from 0.027 – 0.060 Å  and $\Delta\theta_{\rm B-X-B}$ between 12.1 – 20.2. On the right, Cs-based compounds display the largest structural deviations, with $\Delta l_{\rm B-X}$ spanning 0.057 – 0.085 Å  and $\Delta\theta_{\rm B-X-B}$ ranging from 21.1 – 26.5. Notably, MA-based compounds cover the widest range of structural variations, which we attribute to the larger dipole moments of the MA cations that can significantly influence the electronic structure Quarti et al. (2015); Motta et al. (2015).
@@
Figure 5: Dependence of the average B-X-B bond angle decrease on the average B-X bond length variation. The 120 data points correspond to 10 polymorphous configurations generated for 4 FA (black), 4 MA (red), and 4 Cs (blue) -based compounds. The straight thick line represents the least squares regression fit to the data, with the slope indicated on the plot. The shaded region represents three standard deviations on either side of the lines.
@@
Figure 6: (a-f) Convergence of the ASDM anharmonic phonon dispersions at 300 K calculated for MAPbBr3 and FAPbI3 using $2\times 2\times 2$ supercells and the PBEsol approximation. In (a) and (b), we use maximally symmetrized structures (MSS) with space group P422 as the reference. In (c) and (d), we use reference structures with relaxed random molecular (RRM) orientations. In (e) and (f), the molecules adopt a single ferroelectric (FE) orientation. The structures are displayed in the bottom left corner of each plot. The green and red curves represent harmonic phonon dispersions calculated for the monomorphous and polymorphous (ASDM-0) structures, respectively. Subsequent ASDM-$i$ steps, where $i$ represents an integer index for the iteration, are shown as gray and blue. The converged ASDM phonon dispersion is shown as black. (g,h) One-phonon diffuse scattering intensity of MAPbBr3 at 300 K calculated for $Q_{x}$$Q_{y}$0] plane using phonons across the full energy range (g) and phonons with energies below $10$ meV (h). (f) One-phonon diffuse scattering intensity of MAPbBr3 at 300 K calculated for $Q_{x}$$Q_{y}$0.5] plane using phonons with energies below $3.2$ meV. The components of the scattering wavevectors are given in reciprocal lattice units (r.l.u) of the minimal cubic unit-cell.
@@ III.3 Phonon anharmonicity and diffuse scattering
Figure 7: (a,b) Relative free energy at 300 K for 10 ZG polymorphous configurations (r1-r10) of cubic MAPbBr3 (a) and FAPbBr3 (b), referenced to the least stable configuration, calculated using Eq. (8) and $\times2$$\times$2 supercells. The ZG polymorphous configurations are obtained starting from geometry optimization of the reference structures with random molecular orientations and then through ASDM. ASDM converged phonons calculated for each structure [e.g. black lines in Figs. 6(c,d)] are used to compute the free energy. The relative free energy of the ZG polymorphous configuration starting from the maximally symmetrized supercell (MSS) is shown for comparison. (c-f) Band gaps and electron/hole effective masses ($_{\rm e}^{*}/m$$_{\rm h}^{*}$) calculated for MAPbBr3 and FAPbI3 using ZG polymorphous (filled discs) and ZG reference (open squares) configurations in $\times4$$\times$4 supercells accounting for both electron-phonon coupling and thermal lattice expansion. The effect of spin-orbit coupling is also included. Boltzmann-weighted averages [Eq. (11)] are obtained for 10 structures. Open circles represent calculations using the ZG polymorphous configuration of the MSS. All calculated temperature-dependent band gaps are shifted by 1.44 and 1.10 eV to match experimental data of MAPbBr3 and FAPbI3 at 300 K, respectively. These corrections are consistent with the PBE0 values for the band gaps reported in Table 1. Experimental gaps (gray triangles) are from Refs. [Mannino et al., 2020] and [Chen et al., 2020].
To validate our approach for calculating the phonon dispersions of hybrid halide perovskites we report diffuse scattering maps of MAPbBr3 at 300 K. In Fig. 6(g), we show the phonon-induced diffuse scattering map of MAPbBr3 calculated for the $Q_{x}$$Q_{y}$0] plane, where $Q_{\alpha}$ represent Cartesian components of the scattering wavevector $\bf Q$}. The scattering map is computed using the phonons calculated for the MSS structure [Fig. 6(a)], neglecting multi-phonon contributions Zacharias et al. (2021a). Our computed scattering map exhibits excellent qualitative agreement with the experimental maps reported in Ref. [Dubajic et al., 2025], reproducing all main features, including the vertical and horizontal diffuse scattering rods that connect high intensity Bragg peaks, formed when $Q_{x}$ and $Q_{y}$ are integers. In Fig. 6(h), we present the phonon-induced diffuse scattering map of MAPbBr3 in the same plane, considering now only phonons with energies below 10 meV. A comparison of the maps in Figs. 6(g) and (h) clearly reveals that all major features are reproduced, demonstrating that low-energy phonons dominate diffuse scattering in MAPbBr3. In Fig. 6(i), we show the phonon-induced scattering map of MAPbBr3 calculated for the $Q_{x}$$Q_{y}$0.5] plane, focusing now on ultralow energy phonons with energies below 3.2 meV driving dynamical octahedral tilting and hopping between different minima in the PES. These structural fluctuations are primarily associated with the low-energy flat phonon band along the R-M path in Fig. 6(a) and give rise to diffuse scattering rods connecting equivalent wavevectors at the M and R points, consistent with calculations for the high-symmetry cubic phase of CsPbBr3 Zacharias et al. (2023a); Lanigan-Atkins et al. (2021). We note that we exclude any quasielastic contributions to diffuse scattering arising from polymorphism in the inorganic sublattice, as this is beyond the scope of the present work. However, in Sec. V.4, we examine the influence of local disorder on phonons, leading to strongly overdamped vibrational dynamics.
@@ III.4 Free energies, temperature-dependent band gaps, and effective masses
We remark that the 10 polymorphous structures explored for each cubic hybrid halide perovskite in $\times2$$\times$2 supercells yield distinct DFT total energy lowerings and phonon spectra. These variations arise from the sensitivity of local disorder to molecular orientations as discussed previously. To account for the influence of local disorder, configurational entropy, and thermal effects on the electronic structure across various ZG polymorphous configurations, we compute the free energy of each configuration and use it to obtain Boltzmann-weighted averages [Eq. (11)]. Figures 7(a)-(f) show calculations for the relative free energies of each ZG polymorphous configuration and Boltzmann averaged band gaps and effective masses of cubic MAPbBr3 and FAPbI3. The free energies in Figs. 7(a,b) are evaluated with respect to the least stable configurations, r2 and r5. We also include the relative free energy calculated for the ZG polymorphous configuration obtained by combining the ASDM with geometry optimization for the MSS structure. Although starting from the MSS structure yields one of the most stable configurations, it does not lead to the absolute minimum in stability. Other random molecular orientations facilitate local disorder within $\times2$$\times$2 supercells more effectively, leading to lower energy ZG polymorphous structures.
@@
In Figs. 7(d) and (f), we present the temperature dependence of the hole and electron effective masses in cubic MAPbBr3 and FAPbI3, exhibiting a linear relationship with temperature. Compared to calculations for the reference structures, the effective masses obtained using ZG polymorphous configurations are almost twice as large. In fact, for the ZG reference structure of cubic MAPbBr3 we obtain $m^{*}_{\rm h,ref}=0.15$, $m^{*}_{\rm e,ref}=0.14$, and a reduced effective masses of $\mu_{\rm ref}=0.07$ at $T=300$ K in units of the free electron mass $m_{0}$. In contrast, for the ZG polymorphous structure of cubic MAPbBr3, these values increase substantially to $m^{*}_{\rm h,poly}=0.26$, $m^{*}_{\rm e,poly}=0.23$, and $\mu_{\rm poly}=0.12$ at the same temperature. The corresponding values at $T=300$ K for cubic FAPbI3 are $m^{*}_{\rm h,ref}=0.10$, $m^{*}_{\rm e,ref}=0.08$, $\mu_{\rm ref}=0.04$, $m^{*}_{\rm h,poly}=0.16$, $m^{*}_{\rm e,poly}=0.13$, and $\mu_{\rm poly}=0.07$. We note that our calculations are at the DFT-level and can be further improved by considering quasiparticle corrections within the $GW$ approximation, which are known to enhance the effective masses Huang and Lambrecht (2013); Umari et al. (2014); Brivio et al. (2014); Bokdam et al. (2016); Mosconi et al. (2016); Davies et al. (2018); Wang et al. (2020b); Filip et al. (2021); Sajedi et al. (2022) In Sec. IV, we provide a more detailed discussion how the reduced effective masses of polymorphous structures compare to experimental measurements and $GW$ calculations.
Table 1: Lattice constants ($L$), DFT band gaps ($E_{\rm g}$), PBE0 band gaps ($E^{\rm PBE0}_{\rm g}$), DFT hole ($m^{*}_{\rm h}$) and electron ($m^{*}_{\rm e}$) effective masses, and reduced effective masses ($\mu$) of the reference (r) and polymorphous (p) structures of inorganic and hybrid halide perovskites. Band gaps and effective masses are calculated using $2\times 2\times 2$ supercells (40 and 96 atoms), $3\times 3\times 3$ uniform $\bf k$}-grids, and a plane wave cutoff of 120 Ry, including the effect of spin-orbit coupling. DFT corrections to the band gaps $\Delta E_{\rm g}$($T$)] and effective masses $\Delta m^{*}_{\rm h}$($T$)] due to electron-phonon coupling and thermal expansion at temperatures $T$ are calculated with the ASDM using $4\times 4\times 4$ ZG supercells (320 and 768 atoms), $1\times 1\times 1$ uniform $\bf k$}-grids, and a plane wave cutoff of 60 Ry, including the effect of spin-orbit coupling. Values were obtained as averages over 10 distinct configurations of the reference and polymorphous structures. Experimental band gaps of the cubic phases, $E^{\rm expt.}_{\rm g}$, and experimental reduced effective masses of the low-temperature phases ($\mu^{\rm expt.}$) are reported for comparison. $\Delta_{s}$ represents an empirical correction to the calculated band gaps to match $E^{\rm expt.}_{\rm g}$, which is consistent with the PBE0 correction $E^{\rm PBE0}_{\rm g}-E_{\rm g}$.